# iSEARLE: Improving Textual Inversion for Zero-Shot Composed Image Retrieval

### v1

@@ V-C Ablation Studies
We report the results in the lower section of Tab. VIII. The contrastive version of the distillation loss achieves better performance than the cosine one. Compared to the cycle contrastive loss, our distillation-based loss proves to be significantly more effective, highlighting how learning from OTI pre-generated tokens is more fruitful than from raw images. Moreover, although the pre-generated pseudo-word tokens are already regularized, we observe that our GPT-based regularization loss is still beneficial for training $\phi$, especially on the CIRR dataset. Regarding $\mathcal{L}_{pen}$, we find that introducing an additional regularization penalty term to constrain the predicted pseudo-word tokens is effective, as it helps in making them reside in the CLIP manifold [60]. Finally, we observe that by removing the proposed hard negative sampling strategy we achieve comparable performance on CIRR but worse on FashionIQ. This outcome is due to the narrow domain of FashionIQ, as images are closely related and subtle details are crucial. This confirms that our hard negative sampling strategy helps the model in capturing fine-grained details.
TABLE VIII: Ablation studies on CIRR and FashionIQ validation sets. For FashionIQ, we consider the average recall. Best and second-best scores are highlighted in bold and underlined, respectively.
@@
TABLE IX: Evaluation of the effect of the $\phi$ pre-training dataset on the respective evaluation split of FashionIQ, CIRR, and CIRCO. For FashionIQ, we consider the average recall. Best and second-best scores are highlighted in bold and underlined, respectively.
@@ V-D Additional Experiments
Table IX reports the results. We notice how iSEARLE-FIQ achieves the best performance on FashionIQ and even outperforms iSEARLE-OTI, thus highlighting the effectiveness of our distillation-based approach. This result shows that pre-training on images belonging to the same domain as that of the testing ones leads to a performance gain. Moreover, iSEARLE-FIQ also manages to generalize to a broader domain obtaining promising results on both CIRR and CIRCO. Despite relying only on 17K training images, iSEARLE-CIRR achieves noteworthy results, suggesting that our approach is effective even in a low-data regime. Finally, we observe that iSEARLE-FFHQ, iSEARLE-NABirds, and iSEARLE-VGGFace2 obtain promising results despite being pre-trained on datasets that have highly specific domains that are extremely different from those of the testing datasets. To contextualize, Pic2Word [18] scores a $R@1$ of 23.90 on CIRR and a $mAP@10$ of 9.51 on CIRCO. In comparison, iSEARLE-VGGFace2, despite using a smaller backbone and being trained on such a narrow domain as human faces, achieves a $R@1$ of 24.96 on CIRR and a $mAP@10$ of 10.97 on CIRCO. These results show that our approach is robust to the $\phi$ pre-training dataset. Moreover, our model demonstrates noteworthy generalization capabilities, making it well-suited for application in any real-world scenario without the requirement for domain-specific pre-training.
TABLE X: Evaluation of the visual information embedded in $v_{*}$ for different regularization techniques on CIRR validation set. IR and CIR stand for Image Retrieval and Composed Image Retrieval, respectively. Best and second-best scores are highlighted in bold and underlined, respectively.
@@
Visual Information in $\bm{v_{*}}$ We carry out an image retrieval experiment by studying whether the pseudo-word tokens can retrieve the corresponding images. In this way, we assess the effectiveness of the pseudo-word tokens in capturing visual information. Starting from an image $I$, we obtain the corresponding pseudo-word token $v_{*}$ and its associated pseudo-word $S_{*}$ through textual inversion. We craft a generic prompt including the pseudo-word $S_{*}$, such as “a photo of $S_{*}$”. Then, we extract the text features via the CLIP text encoder and use them to query an image database. We expect the image $I$ to be the top-ranked result if the pseudo-word token manages to effectively embed its visual content.
Tab. X shows the results for Image Retrieval (IR) alongside the corresponding ones for Composed Image Retrieval (CIR). We use the CIRR validation set to conduct the experiments. We report the results of all the ablation studies related to the regularization technique for both OTI and $\phi$. Refer to Sec. V-C for more details on the setting of each ablation. Regardless of the regularization strategy, we observe that $v_{*}$ effectively captures the visual information of the image, achieving almost perfect IR scores. However, we obtain a significant performance improvement in CIR when relying on our GPT-powered loss. This underscores how our regularization technique enhances the ability of the pseudo-word tokens to interact with the actual words composing the relative caption while preserving the amount of visual information embedded in $v_{*}$.
@@
TABLE XI: Comparison with supervised baselines on CIRR and FashionIQ validation sets. For FashionIQ, we consider the average recall. Best and second-best scores are highlighted in bold and underlined, respectively.
@@
We measure the generalization capabilities of supervised CIR models by performing a comparison with our zero-shot approach. In particular, we consider Combiner [4], which fuses image and text CLIP features through a combiner network. We chose Combiner as we believe it represents the most similar method to ours among the supervised ones, as we both rely on an out-of-the-box CLIP model. We train two versions of Combiner based on the B/32 backbone on the FashionIQ and CIRR training sets, respectively, using the official repository. We test both Combiner versions on FashionIQ and CIRR validation sets and report the results in Tab. XI. As expected, Combiner achieves the best performance when the training and testing datasets correspond. However, both supervised models struggle to generalize to different domains, as also observed by [18]. Conversely, iSEARLE exhibits remarkable performance on both datasets in a zero-shot manner. Thus, given that we do not require a costly manually annotated training set, the proposed method demonstrates better scalability and suitability for real-world applications of composed image retrieval.

### v2

@@ V-C Ablation Studies
We report the results in the lower section of Tab. VIII. The contrastive version of the distillation loss achieves better performance than the cosine one. Compared to the cycle contrastive loss, our distillation-based loss proves to be significantly more effective, highlighting how learning from OTI pre-generated tokens is more fruitful than learning from raw images. Moreover, although the pre-generated pseudo-word tokens are already regularized, we observe that our GPT-based regularization loss is still beneficial for training $\phi$, especially on the CIRR dataset. Regarding $\mathcal{L}_{pen}$, we find that introducing an additional regularization penalty term to constrain the predicted pseudo-word tokens is effective, as it helps in making them reside in the CLIP manifold [57]. Finally, we observe that by removing the proposed hard negative sampling strategy we achieve comparable performance on CIRR but worse on FashionIQ. This outcome is due to the narrow domain of FashionIQ, as images are closely related and subtle details are crucial. This confirms that our hard negative sampling strategy helps the model in capturing fine-grained details.
Hyperparameters We study the effect of the main hyperparameters of our method and report the results in Tab. IX. Regarding the standard deviation $\gamma$ of the noise, we observe a sweet spot ($\gamma=0.64$) that balances mitigating the modality gap and preserving the informative content of the text features. For the ratio $\alpha$ of hard negative samples per batch, composing batches entirely of hard negatives ($\alpha=1$) leads the model to focus on fine-grained details, improving performance on fine-grained datasets like FashionIQ. However, this comes at the cost of reduced image diversity within each batch, resulting in worse performance on broader-domain datasets such as CIRR. A moderate value ($\alpha=0.5$) achieves a better trade-off by incorporating challenging examples while maintaining a diverse range of image content.
TABLE VIII: Ablation studies on CIRR and FashionIQ validation sets. For FashionIQ, we consider the average recall.
@@
TABLE IX: Ablation studies on the choice of hyperparameters on CIRR and FashionIQ validation sets. For FashionIQ, we consider the average recall.
FashionIQ CIRR
Abl. Value R@10@10@ 10 R@50@50@ 50 R@1@1@ 1 R@5@5@ 5 R@10@10@ 10
$\gamma=0.32$ 24.76 44.73 25.16 56.92 69.98
$\gamma=0.64$ 25.06 44.79 25.57 57.11 70.46
OTI $\gamma=0.96$ 24.15 44.57 25.60 56.54 70.46
$\alpha=0.25$ 23.76 44.49 25.71 57.42 70.28
$\alpha=0.5$ 24.40 44.80 25.74 57.35 70.32
$\phi$ $\alpha=1$ 24.52 44.95 25.14 56.83 70.06
TABLE X: Evaluation of the effect of the $\phi$ pre-training dataset on the respective evaluation split of FashionIQ, CIRR, and CIRCO. For FashionIQ, we consider the average recall.
@@ V-D Additional Experiments
Table X reports the results. We notice how iSEARLE-FIQ achieves the best performance on FashionIQ and even outperforms iSEARLE-OTI, thus highlighting the effectiveness of our distillation-based approach. This result shows that pre-training on images belonging to the same domain as that of the testing ones leads to a performance gain. Moreover, iSEARLE-FIQ also manages to generalize to a broader domain obtaining promising results on both CIRR and CIRCO. Despite relying only on 17K training images, iSEARLE-CIRR achieves noteworthy results, suggesting that our approach is effective even in a low-data regime. Finally, we observe that iSEARLE-FFHQ, iSEARLE-NABirds, and iSEARLE-VGGFace2 obtain promising results despite being pre-trained on datasets that have highly specific domains that are extremely different from those of the testing datasets. To contextualize, Pic2Word [18] scores a $R@1$ of 23.90 on CIRR and a $mAP@10$ of 9.51 on CIRCO. In comparison, iSEARLE-VGGFace2, despite using a smaller backbone and being trained on such a narrow domain as human faces, achieves a $R@1$ of 24.96 on CIRR and a $mAP@10$ of 10.97 on CIRCO. These results show that our approach is robust to the $\phi$ pre-training dataset. Moreover, our model demonstrates noteworthy generalization capabilities, making it well-suited for application in any real-world scenario without the requirement for domain-specific pre-training.
TABLE XI: Evaluation of the visual information embedded in $v_{*}$ for different regularization techniques on CIRR validation set. IR and CIR stand for Image Retrieval and Composed Image Retrieval, respectively.
@@
Visual Information in $\bm{v_{*}}$ We carry out an image retrieval experiment by studying whether the pseudo-word tokens can retrieve the corresponding images. In this way, we assess the effectiveness of the pseudo-word tokens in capturing visual information. Starting from an image $I$, we obtain the corresponding pseudo-word token $v_{*}$ and its associated pseudo-word $S_{*}$ through textual inversion. We craft a generic prompt including the pseudo-word $S_{*}$, such as ``a photo of $S_{*}$". Then, we extract the text features via the CLIP text encoder and use them to query an image database. We expect the image $I$ to be the top-ranked result if the pseudo-word token manages to effectively embed its visual content.
Tab. XI shows the results for Image Retrieval (IR) alongside the corresponding ones for Composed Image Retrieval (CIR). We use the CIRR validation set to conduct the experiments. We report the results of all the ablation studies related to the regularization technique for both OTI and $\phi$. Refer to Sec. V-C for more details on the setting of each ablation. Regardless of the regularization strategy, we observe that $v_{*}$ effectively captures the visual information of the image, achieving almost perfect IR scores. However, we obtain a significant performance improvement in CIR when relying on our GPT-powered loss. This highlights how our regularization technique enhances the ability of the pseudo-word tokens to interact with the actual words composing the relative caption while preserving the visual information embedded in $v_{*}$.
@@
TABLE XII: Comparison with supervised baselines on CIRR and FashionIQ validation sets. For FashionIQ, we consider the average recall.
@@
We measure the generalization capabilities of supervised CIR models by performing a comparison with our zero-shot approach. In particular, we consider Combiner [4], which fuses image and text CLIP features through a combiner network. We chose Combiner as we believe it represents the most similar method to ours among the supervised ones, as we both rely on an out-of-the-box CLIP model. We train two versions of Combiner based on the B/32 backbone on the FashionIQ and CIRR training sets, respectively, using the official repository. We test both Combiner versions on FashionIQ and CIRR validation sets and report the results in Tab. XII. As expected, Combiner achieves the best performance when the training and testing datasets correspond. However, both supervised models struggle to generalize to different domains, as also observed by [18]. Conversely, iSEARLE exhibits remarkable performance on both datasets in a zero-shot manner. Thus, given that we do not require a costly manually annotated training set, the proposed method demonstrates better scalability and suitability for real-world applications of composed image retrieval.